Circular No. 7710

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ef IN IC 381 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong (Rolvenden, England) of an apparent supernova (mag 16.1) on unfiltered CCD images taken on Sept. 9.066 and 9.160 UT in the course of the U.K. Nova/Supernova Patrol with a 0.36-m Schmidt-Cassegrain telescope. An image by Armstrong on Sept. 9.907 yields mag 16.0; a confirming image was also obtained by N. James (Chelmsford, England) on Sept. 9.900. SN 2001ef is located at α = 4^{h}44^{m}28^{s}.56, δ = +75^{o}38′30′′.3 (equinox 2000.0; average of three measures), which is 1′′ east and 8′′ north of the center of IC 381. Nothing appears at this location on images taken by Armstrong on 2000 Oct. 9 (limiting mag 19.5) and 2001 Aug. 17 and 29; nothing is visible on second Palomar Sky Survey images taken on 1995 May 19 (blue, limiting mag 22.5), 1996 Dec. 8 (red, limiting mag 20.8), and 1997 July 19 (infrared). Hurst adds that nothing is present at this position in an image in J. C. Vickers’ Deep Space CCD Atlas North (1993, p. 44; limiting mag ∼ 19.5). A. Miceli, C. Stubbs, and C. Hastings find the new object to be 0′′.33 east and 7′′.3 north of the galaxy’s center; their preliminary photometry with the Apache Point 3.5-m telescope on Sept. 10.44 yields u′ ∼ 17.5, g′ ∼ 17.3, r′ ∼ 16.5, i′ ∼ 16.9, z′ ∼ 16.1.

XTE J1650–500 T. Augusteijn, Isaac Newton Group of Telescopes, La Palma; M. Coe,

University of Southampton; and P. Groot, Harvard-Smithsonian Center for Astrophysics, report: “Spectroscopic observations (range 390–760 nm; resolution 0.7 nm) of the proposed optical counterpart of the x-ray transient XTE J1650–500 (cf. IAUC 7707, 7708) were made on Sept. 8.74 UT with the 1.9-m telescope at the South African Astronomical Observatory. The observations were made through varying cloud cover, and are not flux-calibrated. The spectrum shows emission lines from the Balmer series, which are unresolved and have equivalent width (EW) 0.2–0.4 nm. This confirms the identification of the source as an (soft-x-ray-type) low-mass x-ray binary in outburst. Also detected are several interstellar absorption features. From the EW of 0.20 ± 0.02 nm obtained for the diffuse interstellar band at 628.3 nm, we estimate an E(B − V) ∼ 1.5 mag. Assuming that it is at the distance of the Galactic center, we derive an absolute visual magnitude of −2 for the source. This implies that the source is likely a long-orbital-period system, as is also indicated by the apparent small outburst amplitude.”

2001 September 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green